# On functoriality and the tensor product property in noncommutative tensor-triangular geometry

### v2

@@ Abstract.
Two pertinent questions for any support theory of a monoidal triangulated category are whether it is functorial and if the tensor product property holds. To this end, we consider the complete prime spectrum $\operatorname{Spc}^{cp}(\mathscr{K})$ of an essentially small monoidal triangulated category $\mathscr{K}$, which we show is universal among support data on $\mathscr{K}$ satisfying the tensor product property, even if it is empty. We show $\operatorname{Spc}^{cp}(-)$ is functorial and give a characterization of when a monoidal exact functor $F$ induces a well-defined map on noncommutative Balmer spectra. We give criteria for when induced maps on complete prime spectra are injective or surjective, and determine the complete prime spectrum for crossed product categories. Finally, we determine the universal functorial support theory for monoidal triangulated categories coinciding with the Balmer spectrum on braided monoidal triangulated categories.
@@ 1. Introduction
Theorem (Theorem 8.5).
Let $\mathscr{K}=D_{b}(k)\oplus D_{b}(k)$. The crossed product category $\mathscr{K}\rtimes C_{2}$, with $C_{2}$ acting by permuting indices, contains no completely prime ideals. Moreover, the inclusion $F:\mathscr{K}\hookrightarrow\mathscr{K}\rtimes C_{2}$ does not induce a well-defined map $\operatorname{Spc}(F):\operatorname{Spc}(\mathscr{K}\rtimes C_{2})\to\operatorname{Spc}(\mathscr{K})$ nor a well-defined functor (see Definition 3.1) $\mathscr{S}(F):\mathscr{S}(\mathscr{K}\rtimes C_{2})\to\mathscr{S}(\mathscr{K})$.
More generally, we give a criteria for precisely when a monoidal exact functor $F:\mathscr{K}\to\mathscr{L}$ induces a well-defined map on Balmer spectra. As a result, we can deduce the noncommutative Balmer spectra of localizations and idempotent completion, adaptations of standard tensor-triangular geometric results for the noncommutative setting.
Theorem (Theorem 4.4).
@@
If we think of this complete prime spectrum $\operatorname{Spc}^{cp}(\mathscr{K})$ as another construction of a spectral space or support theory, then in some ways it is better-behaved; for instance, it is indeed functorial. However, it has its own major shortcomings (as explained in [NVY22a]) - notably, it may be empty, and the standard “prime lifting” techniques (see [Bal05, Lemma 2.2] or [NVY22a, Theorem 3.2.3]) do not seem feasible if $\operatorname{Spc}^{cp}(\mathscr{K})\subset\operatorname{Spc}(\mathscr{K})$. Frame-theoretically, it is insufficient. However, in tensor-triangular geometry, one should think about the Balmer spectrum in the terms of universal support: the (commutative) Balmer spectrum is the universal support datum for tensor-triangulated categories, and the noncommutative Balmer spectrum is universal for, in fact, multiple notions of noncommutative support data. And similarly, the complete prime spectrum is itself universal - it is the final object in the category of multiplicative support data, i.e. support data satisfying the tensor product property.
Theorem.
@@
  • (Theorem 3.2) $\mathscr{S}^{\otimes}(-):\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Cat}$ is a contravariant functor from the category of essentially small monoidal triangulated categories to the category of essentially small categories.
  • (Theorem 5.2) The complete prime spectrum $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$ is the final object of $\mathscr{S}^{\otimes}(\mathscr{K})$, even if $\operatorname{Spc}^{cp}(\mathscr{K})=\emptyset$.
  • (Theorem 5.5) The assignment $\operatorname{Spc}^{cp}(-):\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ is a contravariant functor from the category of essentially small monoidal triangulated categories to the category of topological spaces.
Given a spectral theory on a monoidal triangulated category $\mathscr{K}$, the specialization-closed or Thomason subsets should be expected to parametrize some data about $\mathscr{K}$. We show that with some modifications, $\operatorname{Spc}^{cp}(\mathscr{K})$ parametrizes certain thick tensor-ideals.
Theorem (Theorem 6.8).
The set of “dual-open” subsets of $\operatorname{Spc}^{cp}(\mathscr{K})$, i.e. arbitrary unions of supports $\operatorname{supp}(a)$ for $a\in\mathscr{K}$, is in order-preserving bijection with the set of radical thick tensor-ideals of $\mathscr{K}$ that are intersections of complete primes.
If $\operatorname{Spc}^{cp}(\mathscr{K})=\operatorname{Spc}(\mathscr{K})$, then one may replace “dual-open” with Thomason subsets and remove the final intersection condition, in other words, we recover [Bal05, Theorem 4.10]. If $\operatorname{Spc}^{cp}(\mathscr{K})$ is noetherian, one may replace “dual-open” with specialization closed.
@@
  • (Corollary 11.12) The functor $p\operatorname{Spc}:\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ is the final object in the full subcategory of $\operatorname{Fun}(\mathbf{mon}_{\Delta}^{\operatorname{op}},\mathbf{Top})$ consisting of functors $F:\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ which are naturally isomorphic to $\operatorname{Spc}$ upon restriction to $\mathbf{Bmon}_{\Delta}^{\operatorname{op}}$.
  • (Theorem 11.15) If $p\operatorname{Spc}(\mathscr{K})\neq\emptyset$, then the pair $(p\operatorname{Spc}(\mathscr{K}),p\operatorname{supp})$ is a final object in the category $\mathscr{S}^{cm}(\mathscr{K})$ of commutative support data (see Definition 11.14).
@@ 2. Preliminaries
$\mathscr{I}\otimes\mathscr{J}\subseteq\mathscr{P}\implies\mathscr{I}\subseteq\mathscr{P}\text{ or }\mathscr{J}\subseteq\mathscr{P}$
@@
$x\otimes y\in\mathscr{P}\implies x\in\mathscr{P}\text{ or }y\in\mathscr{P}.$
@@ 3. Categories of support data
  • Given an exact $\otimes$-functor $F:\mathscr{K}\to\mathscr{L}$, when is the induced map on Balmer spectra injective or surjective?
@@ 4. Partial functoriality of the noncommutative Balmer spectrum
Why is functoriality important, besides simply being a nice property to have? We turn to the classical symmetric monoidal category case for inspiration. Often when the (commutative) Balmer spectrum cannot be discerned by standard means such as comparison maps as in [Bal10a] or [NVY24], it is discerned from “fiber functors,” a family of functors which together detect vanishing or tensor-nilpotence of morphisms (certain techniques allow for weaker assumptions than conservativity, see [BCHS24, Definition 1.1]). For instance, Balmer and Gallauer in [BG23] construct modular fixed points functors to recover $\operatorname{Spc}(\operatorname{K}_{b}(p\operatorname{-perm}(kG)))$ as a set, where $\operatorname{K}_{b}(p\operatorname{-perm}(kG))$ denotes the bounded homotopy category of $p$-permutation $kG$-modules. Similarly, in [BS17], Balmer and Sanders use geometric fixed points functors to deduce $\operatorname{Spc}(\operatorname{SH}(G)^{c})$, where $\operatorname{SH}(G)^{c}$ denotes the compact part of the $G$-equivariant stable homotopy category. Conservativity techniques have been developed in detail for symmetric monoidal categories in [Bal18], [BCHS24] and recently [BG25].
@@
$\operatorname{supp}_{\mathscr{L}}(F(x))$ =$=\{\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})\mid F(x)\not\in\mathscr{Q}\}$
@@
We can in fact give a criterion for when an exact $\otimes$-functor $F:\mathscr{K}\to\mathscr{L}$ induces a well-defined map on noncommutative Balmer spectra.
@@
It is routine to verify that for every thick $\otimes$-ideal $\mathscr{J}\subseteq\mathscr{L}$, $F^{-1}(\mathscr{J})$ is a thick $\otimes$-ideal of $\mathscr{K}$. Therefore, for $\operatorname{Spc}(F)$ is well-defined if and only if the primality property is preserved for all primes $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$. We have that $F$, as a monoidal (but not necessarily triangulated, since $F(\mathscr{K})$ may not be triangulated if $F$ is not full) functor factors as the surjective monoidal functor $F:\mathscr{K}\to F(\mathscr{K})$ composed with the inclusion $\iota:F(\mathscr{K})\subseteq\mathscr{L}$. The same argument as Proposition 4.2 demonstrates that for every $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, $F^{-1}(F(\mathscr{K})\cap\mathscr{Q})\subseteq\mathscr{K}$ satisfies the tensor product property by surjectivity of $F:\mathscr{K}\to F(\mathscr{K}).$ We turn to the inclusion $\iota$.
@@
$\widetilde{\mathscr{P}}:=\{x\in\mathscr{L}\mid x\oplus\Sigma x\in\mathscr{P}\}=\{x\in\mathscr{L}\mid\exists x^{\prime}\in\mathscr{L}\text{ such that }x\oplus x^{\prime}\in\mathscr{P}\}.$
@@ 5. Functoriality and universality of the complete prime spectrum
Let $F:\mathscr{K}\to\mathscr{L}$ be a (necessarily nonzero) exact $\otimes$-functor and suppose $\operatorname{Spc}^{cp}(\mathscr{L})\neq\emptyset$. Then $\operatorname{Spc}^{cp}(\mathscr{K})\neq\emptyset$.
@@ 6. Radical thick tensor-ideals
In Remark 3.5, we claimed that any reasonable choice of support theory axioms should satisfy that the specialization-closed subsets of its Balmer spectrum should parametrize something about our category $\mathscr{K}$. We’ll investigate the specialization closed subsets for multiplicative support data, i.e. the complete prime spectrum $\operatorname{Spc}^{cp}(\mathscr{K})$. Perhaps unsurprisingly, this verification mirrors the classical setting of [Bal05], where radical thick $\otimes$-ideals come into play.
@@
Let $\mathscr{I}$ be a thick $\otimes$-ideal of $\mathscr{K}$. We say $\mathscr{I}$ is radical if $\sqrt{\mathscr{I}}=\mathscr{I}$ or equivalently, for all $x\in\mathscr{K}$, $x\in\mathscr{I}$ if and only if $x^{\otimes n}\in\mathscr{I}$ for all $n\geq 1$.
For symmetric tensor categories, every radical thick $\otimes$-ideal equivalently the intersection of primes, see [Bal05, Lemma 4.2]. In our setting, we have a weaker version.
Proposition 6.2.
Every intersection of completely prime thick $\otimes$-ideals is a radical $\otimes$-ideal. Conversely, every radical thick $\otimes$-ideal $\mathscr{I}$ is the intersection of some family of prime $\otimes$-ideals. In particular, we have
$\mathscr{I}=\bigcup_{\mathscr{P}\supseteq\mathscr{I}}\mathscr{P}.$
Proof.
First, let $\mathscr{I}$ be the intersection of some family of completely prime thick $\otimes$-ideals $\{\mathscr{P}_{i}\}_{i\in I}$. Suppose $x^{\otimes n}\in\mathscr{I}$. Then for every $\mathscr{P}_{i}$, $x^{\otimes n}\in\mathscr{P}_{i}$, and complete primality implies $x\in\mathscr{P}_{i}$, whence $x\in\mathscr{I}$.
Now, suppose $\mathscr{I}$ is a radical thick prime $\otimes$-ideal. We check the proposed equality; it is clear that $\mathscr{I}\subseteq\bigcup_{\mathscr{P}\supseteq\mathscr{I}}\mathscr{P}$. Conversely, suppose $x\in\mathscr{P}$ for all $\mathscr{P}\in\operatorname{Spc}(\mathscr{K})$ containing $\mathscr{I}$, and suppose for contradiction that $x\not\in\mathscr{I}$. Then we have a multiplicative set $S=\{x,x^{\otimes 2},\dots\}$ intersecting $\mathscr{I}$ trivially, and prime lifting 2.4 asserts the existence of a prime containing $\mathscr{I}$ and not containing $x$, a contradiction. ∎
Remark 6.3.
A pertinent question is whether every radical thick $\otimes$-ideal of $\mathscr{K}$ is an intersection of completely prime thick $\otimes$-ideals. At present, we are not aware of a counterexample.
On the other hand, an intersection of only prime thick $\otimes$-ideals is not necessarily radical, since not all prime thick $\otimes$-ideals are radical. In fact, for prime thick $\otimes$-ideals, being radical is equivalent to being completely prime.
@@
Proposition 6.5.
Given any specialization closed subset $Y\subseteq\operatorname{Spc}^{cp}(\mathscr{K})$, one may construct the full subcategory of $\mathscr{K}$
$\mathscr{K}^{cp}_{Y}:=\{x\in\mathscr{K}\mid\operatorname{supp}^{cp}(x)\subseteq Y\}\subseteq\mathscr{K}.$
We have that $\mathscr{K}^{cp}_{Y}$ is a radical thick $\otimes$-ideal of $\mathscr{K}$. Equivalently, $\mathscr{K}^{cp}_{Y}$ is equal to the intersection $\bigcap_{\mathscr{P}\not\in Y}\mathscr{P}$ of all complete primes $\mathscr{P}\in\operatorname{Spc}^{cp}(\mathscr{K})$ not belonging to $Y$.
@@
The verification that $\mathscr{K}^{cp}_{Y}$ is a thick $\otimes$-ideal is straightforward and analogous to [NVY22a, Proposition 6.1.1]. Because $\operatorname{supp}^{cp}(x^{\otimes n})=\bigcap_{n}\operatorname{supp}^{cp}(x)=\operatorname{supp}^{cp}(x)$ for all $x\in\mathscr{K}$ and $n\geq 1$, $\mathscr{K}^{cp}_{Y}$ is radical. The last statement is just [Bal05, Lemma 4.8]; its proof is purely set-theoretic and follows the same here. ∎
Definition 6.6.
Given a collection of objects $S\subseteq\mathscr{K}$, we set
$\operatorname{supp}^{cp}(\mathscr{I}):=\bigcup_{x\in S}\operatorname{supp}^{cp}(x).$
Since $\operatorname{supp}(\mathscr{I})$ is a union of supports, it is specialization-closed. Moreover, observe $\operatorname{supp}(S)=\{\mathscr{P}\in\operatorname{Spc}^{cp}(\mathscr{K})\mid S\not\subseteq\mathscr{P}\}$.
We now come to our classification theorem, in the spirit of [Bal05, Theorem 4.10] and [NVY22a, Theorem 6.2.1]. It is certainly a less satisfying statement than the aforementioned theorems, but sadly it seems the complete prime spectrum is not as well-suited for classification. There is a technical point - in this setting it is not clear if every open $U^{cp}(x)$ is quasi-compact, (although this holds if every prime is complete). Therefore, we must replace Thomason subsets, i.e. arbitrary unions of closed subsets with quasi-compact compliment, with the more explicit “arbitrary unions of supports,” i.e. dual-opens. However, if $\operatorname{Spc}^{cp}(\mathscr{K})$ is noetherian, we may replace this condition with specialization closed, since every arbitrary union of closed subsets is in fact a finite union.
In general, it is unclear if $U^{cp}(a)\subseteq\operatorname{Spc}^{cp}(\mathscr{K})$ is always quasi-compact, though we know every quasi-compact open is of the form $U^{cp}(a)$. However, this holds when every prime is completely prime. The proof of this fact follows verbatim.
Proposition 6.7.
[Bal05, Proposition 2.14] Assume $\operatorname{Spc}(\mathscr{K})=\operatorname{Spc}^{cp}(\mathscr{K})$. The following facts hold:
  1. (a)
    For any object $x\in\mathscr{K}$, the open $U(x)$ is quasi-compact.
  2. (b)
    Conversely, any quasi-compact open of $\operatorname{Spc}(\mathscr{K})$ is of the form $U(x)$ for some $x\in\mathscr{K}$.
In particular, $\operatorname{Spc}(\mathscr{K})$ is noetherian if and only if any closed subset of $\operatorname{Spc}(\mathscr{K})$ is the support of an object of $\mathscr{K}$.
Theorem 6.8.
Let $\mathfrak{S}$ be the set of dual-open subsets of $\operatorname{Spc}^{cp}(\mathscr{K})$, and let $\mathfrak{R}$ be the set of radical thick $\otimes$-ideals of $\mathscr{K}$ which are an intersection of complete prime thick $\otimes$-ideals. Then $\mathfrak{S}$ and $\mathfrak{R}$ are in order-preserving bijection via the assignment
$Y\mapsto\mathscr{K}_{Y}^{cp}=\{x\in\mathscr{K}\mid\operatorname{supp}^{cp}(x)\subseteq Y\}$
with inverse
$\mathscr{I}\mapsto\operatorname{supp}^{cp}(\mathscr{I})=\bigcup_{x\in\mathscr{I}}\operatorname{supp}^{cp}(x).$
Proof.
The well-definedness and order preservation of the two assignments is clear from construction; it suffices to verify that they are indeed inverse. First, we consider $\mathscr{K}_{\operatorname{supp}^{cp}(\mathscr{I})}^{cp}$. By Proposition 6.5, we have $\mathscr{K}_{\operatorname{supp}^{cp}(\mathscr{I})}^{cp}=\bigcap_{\mathscr{P}\not\in\operatorname{supp}^{cp}(\mathscr{I})}\mathscr{P}$, and equivalently $\mathscr{P}\not\in\operatorname{supp}^{cp}(\mathscr{I})$ if and only if $\mathscr{I}\subseteq\mathscr{P}$, so
$\mathscr{K}^{cp}_{\operatorname{supp}^{cp}(\mathscr{I})}=\bigcap_{\mathscr{I}\subseteq\mathscr{P},\mathscr{P}\in\operatorname{Spc}^{cp}(\mathscr{K})}\mathscr{P}.$
Since we specified $\mathscr{I}$ is explicitly an intersection of complete primes and radical, we conclude $\mathscr{I}=\mathscr{K}^{cp}_{\operatorname{supp}^{cp}(\mathscr{I})}$, since $\mathscr{K}^{cp}_{\operatorname{supp}^{cp}(\mathscr{I})}$ is the smallest intersection of complete primes containing $\mathscr{I}$, which must be $\mathscr{I}$ itself.
Now, consider $\operatorname{supp}^{cp}(\mathscr{K}_{Y})$ for some intersection of supports $Y\subseteq\operatorname{Spc}^{cp}(\mathscr{K})$. First if $\mathscr{P}\in\operatorname{supp}^{cp}(\mathscr{K}_{Y})$, then $\mathscr{P}\in\operatorname{supp}^{cp}(x)$ for some $x\in\mathscr{K}_{Y}$. By definition of $\mathscr{K}^{cp}_{Y}$, $x$ satisfies that $\operatorname{supp}^{cp}(x)\subseteq Y$, so $\mathscr{P}\in Y$. Conversely, suppose $\mathscr{P}\in Y$. Since we have assumed $Y$ is an arbitrary union of supports, if $\mathscr{P}\in Y$, then $\mathscr{P}\in\operatorname{supp}^{cp}(x)\subseteq Y$ for some $x\in\mathscr{K}$. By definition, $x\in\mathscr{K}^{cp}_{Y}$, so therefore $\mathscr{P}\in\bigcup_{x\in\mathscr{K}^{cp}_{Y}}\operatorname{supp}(x)=\operatorname{supp}(\mathscr{K}^{cp}_{Y})$, and we conclude the assignments are indeed inverse. ∎
Corollary 6.9.
If $\operatorname{Spc}(\mathscr{K})=\operatorname{Spc}^{cp}(\mathscr{K})$, then the bijection of Theorem 6.8 is an order-preserving bijection between the Thomason-closed subsets of $\operatorname{Spc}(\mathscr{K})$ and the radical thick $\otimes$-ideals of $\mathscr{K}$. If $\operatorname{Spc}^{cp}(\mathscr{K})$ is a noetherian topological space, then the bijection is an order-preserving bijection between specialization-closed subsets and radical thick $\otimes$-ideals of $\mathscr{K}$ which are an intersection of complete primes.
@@
We note that one can say more about the noncommutative Balmer spectrum when $\operatorname{Spc}(\mathscr{K})=\operatorname{Spc}^{cp}(\mathscr{K})$: nearly every topological property proven in [Bal05, Section 2] holds, essentially due to the prime lifting lemma Proposition 2.4 holding for complete primes.

### v3

@@ Abstract.
Two pertinent questions for any support theory of a monoidal triangulated category are whether it is functorial and if the tensor product property holds. To this end, we consider the complete prime spectrum of an essentially small monoidal triangulated category, which we show is universal among support data satisfying the tensor product property, even if it is empty. The complete prime spectrum is functorial and parametrizes radical thick tensor-ideals, a noncommutative analogue of Balmer’s reconstruction theorem. We give criteria for when induced maps on complete prime spectra are injective or surjective, and determine the complete prime spectrum for crossed product categories. Finally, we determine the universal functorial support theory for monoidal triangulated categories coinciding with the Balmer spectrum on braided monoidal triangulated categories.
@@ 1. Introduction
Theorem A (Theorem 8.5).
Let $\mathscr{K}=D_{b}(k)\oplus D_{b}(k)$. The crossed product category $\mathscr{K}\rtimes C_{2}$, with $C_{2}$ acting by permuting indices, contains no completely prime ideals. Moreover, the inclusion $F\colon\mathscr{K}\hookrightarrow\mathscr{K}\rtimes C_{2}$ does not induce a well-defined map $\operatorname{Spc}(F):\operatorname{Spc}(\mathscr{K}\rtimes C_{2})\to\operatorname{Spc}(\mathscr{K})$ nor a well-defined functor (see Definition 3.1) $\mathscr{S}(F):\mathscr{S}(\mathscr{K}\rtimes C_{2})\to\mathscr{S}(\mathscr{K})$.
More generally, we give a criteria for precisely when a monoidal exact functor $F\colon\mathscr{K}\to\mathscr{L}$ induces a well-defined map on Balmer spectra. As a result, we can deduce the noncommutative Balmer spectra of localizations and idempotent completion, adaptations of standard tensor-triangular geometric results for the noncommutative setting.
Theorem B (Theorem 4.4).
@@
If we think of this complete prime spectrum $\operatorname{Spc}^{cp}(\mathscr{K})$ as another construction of a spectral space or support theory, then in some ways it is better-behaved; for instance, it is indeed functorial. However, it has its own major shortcomings (as explained in [NVY22a]) - notably, it may be empty, and the standard “prime lifting” techniques (see [Bal05, Lemma 2.2] or [NVY22a, Theorem 3.2.3]) do not seem feasible if $\operatorname{Spc}^{cp}(\mathscr{K})\subset\operatorname{Spc}(\mathscr{K})$. However, in tensor-triangular geometry, one should think about the Balmer spectrum in the terms of universal support: the (commutative) Balmer spectrum is the universal support datum for tensor-triangulated categories, and the noncommutative Balmer spectrum is universal for, in fact, multiple notions of noncommutative support data. And similarly, the complete prime spectrum is itself universal - it is the final object in the category of multiplicative support data, i.e. support data satisfying the tensor product property.
Theorem C.
@@
  • $\mathscr{S}^{\otimes}(-):\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Cat}$ is a contravariant functor from the category of essentially small monoidal triangulated categories to the category of essentially small categories; (Theorem 3.2)
  • The complete prime spectrum $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$ is the final object of $\mathscr{S}^{\otimes}(\mathscr{K})$, even if $\operatorname{Spc}^{cp}(\mathscr{K})=\emptyset$; (Theorem 5.2)
  • The assignment $\operatorname{Spc}^{cp}(-):\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ is a contravariant functor from the category of essentially small monoidal triangulated categories to the category of topological spaces. (Theorem 5.5)
Given a spectral theory on a monoidal triangulated category $\mathscr{K}$, the specialization-closed or Thomason subsets should be expected to parametrize some data about $\mathscr{K}$. Reframing slightly, we prove an analogue of Balmer’s reconstruction theorem [Bal05, Theorem 4.10]. As in the commutative setting, the complete primes parametrize the radical thick $\otimes$-ideals of $\mathscr{K}$. In fact, a complete prime is nothing more than a radical prime.
Theorem D.
Let $\operatorname{Spc}^{cp}(\mathscr{K})^{\vee}$ denote the Hochster dual of $\operatorname{Spc}^{cp}(\mathscr{K})$ (see Definition 6.3).
  1. (a)
    Let $\mathscr{P}$ be a thick $\otimes$-ideal of $\mathscr{K}$. Then $\mathscr{P}$ is completely prime if and only if $\mathscr{P}$ is prime and radical; (Proposition 6.2)
  2. (b)
    We have an order-preserving bijection from the set $\mathscr{R}$ of radical thick $\otimes$-ideals to the set $\mathscr{S}$ of open subsets of $\operatorname{Spc}^{cp}(\mathscr{K})^{\vee}$ induced by the assignment
    $\mathscr{I}\mapsto\bigcup_{x\in\mathscr{I}}\operatorname{supp}(x).$
    In particular, every radical thick $\otimes$-ideal of $\mathscr{K}$ is an intersection of completely prime thick $\otimes$-ideals of $\mathscr{K}$. (Theorem 6.4)
@@
  • The functor $p\operatorname{Spc}:\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ is the final object in the full subcategory of $\operatorname{Fun}(\mathbf{mon}_{\Delta}^{\operatorname{op}},\mathbf{Top})$ consisting of functors $F\colon\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ which are naturally isomorphic to $\operatorname{Spc}$ upon restriction to $\mathbf{Bmon}_{\Delta}^{\operatorname{op}}$; (Corollary 11.12)
  • If $p\operatorname{Spc}(\mathscr{K})\neq\emptyset$, then the pair $(p\operatorname{Spc}(\mathscr{K}),p\operatorname{supp})$ is a final object in the category $\mathscr{S}^{cm}(\mathscr{K})$ of commutative support data (see Definition 11.14). (Theorem 11.15)
@@ 2. Preliminaries
[…]
@@
[…]
@@ 3. Categories of support data
  • Given an exact $\otimes$-functor $F\colon\mathscr{K}\to\mathscr{L}$, when is the induced map on Balmer spectra injective or surjective?
@@ 4. Partial functoriality of the noncommutative Balmer spectrum
Why is functoriality important, besides simply being a nice property to have? We turn to the classical symmetric monoidal category case for inspiration. Often when the (commutative) Balmer spectrum cannot be discerned by standard means such as comparison maps as in [Bal10a] or [NVY24], it is discerned from “fiber functors,” a family of functors which together detect vanishing or tensor-nilpotence of morphisms (certain techniques allow for weaker assumptions than conservativity, see [BCHS24, Definition 1.1]). For instance, Balmer and Gallauer in [BG23] construct modular fixed points functors to recover $\operatorname{Spc}(\operatorname{K}_{b}(p\operatorname{-perm}(kG)))$ as a set, where $\operatorname{K}_{b}(p\operatorname{-perm}(kG))$ denotes the bounded homotopy category of $p$-permutation kGkG-modules. Similarly, in [BS17], Balmer and Sanders use geometric fixed points functors to deduce $\operatorname{Spc}(\operatorname{SH}(G)^{c})$, where $\operatorname{SH}(G)^{c}$ denotes the compact part of the $G$-equivariant stable homotopy category. Conservativity techniques have been developed in detail for symmetric monoidal categories in [Bal18], [BCHS24] and recently [BG25].
@@
$\operatorname{supp}_{\mathscr{L}}(F(x))$ $=\{\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})\mid F(x)\not\in\mathscr{Q}\}$
@@
We can in fact give a criterion for when an exact $\otimes$-functor $F\colon\mathscr{K}\to\mathscr{L}$ induces a well-defined map on noncommutative Balmer spectra.
@@
It is routine to verify that for every thick $\otimes$-ideal $\mathscr{J}\subseteq\mathscr{L}$, $F^{-1}(\mathscr{J})$ is a thick $\otimes$-ideal of $\mathscr{K}$. Therefore, for $\operatorname{Spc}(F)$ is well-defined if and only if the primality property is preserved for all primes $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$. We have that $F$, as a monoidal (but not necessarily triangulated, since $F(\mathscr{K})$ may not be triangulated if $F$ is not full) functor factors as the surjective monoidal functor $F\colon\mathscr{K}\to F(\mathscr{K})$ composed with the inclusion $\iota:F(\mathscr{K})\subseteq\mathscr{L}$. The same argument as Proposition 4.2 demonstrates that for every $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, $F^{-1}(F(\mathscr{K})\cap\mathscr{Q})\subseteq\mathscr{K}$ satisfies the tensor product property by surjectivity of $F\colon\mathscr{K}\to F(\mathscr{K}).$ We turn to the inclusion $\iota$.
@@
[…]
@@ 5. Functoriality and universality of the complete prime spectrum
Let $F\colon\mathscr{K}\to\mathscr{L}$ be a (necessarily nonzero) exact $\otimes$-functor and suppose $\operatorname{Spc}^{cp}(\mathscr{L})\neq\emptyset$. Then $\operatorname{Spc}^{cp}(\mathscr{K})\neq\emptyset$.
@@ 6. Radical thick tensor-ideals
In Remark 3.5, we claimed that any reasonable choice of support theory axioms should satisfy that the specialization-closed subsets of its Balmer spectrum should parametrize something about our category $\mathscr{K}$. Indeed, perhaps unsurprisingly we will see in this section that $\operatorname{Spc}^{cp}(\mathscr{K})$ parametrizes radical thick $\otimes$-ideals.
@@
Let $\mathscr{I}$ be a thick $\otimes$-ideal of $\mathscr{K}$. We say $\mathscr{I}$ is radical if $\sqrt{\mathscr{I}}=\mathscr{I}$ or equivalently, for all $\in\mathscr{K}, xx\in\mathscr{I}$ if and only if $x^{\otimes n}\in\mathscr{I}$ for all $n\geq 1$.
In fact, complete primes are nothing more than radical primes.
@@
Now, following [Kra24], we have an analogue of [Bal05, Theorem 4.10]. However we must be careful - rather than using the language of Thomason subsets, we instead utilize the Hochster dual topology on $\operatorname{Spc}^{cp}(\mathscr{K})$.
Definition 6.3.
We define the Hochster dual topology of $\operatorname{Spc}^{cp}(\mathscr{K})$, denoted
$\operatorname{Spc}^{cp}(\mathscr{K})^{\vee}$, as follows. As a set, $\operatorname{Spc}^{cp}(\mathscr{K})^{\vee}:=\operatorname{Spc}^{cp}(\mathscr{K})$, and the supports $\operatorname{supp}(x)$ for $x\in\mathscr{K}$ generate yield a basis of open sets for the topology of $\operatorname{Spc}^{cp}(\mathscr{K})^{\vee}$.
Theorem 6.4.
We have an order-preserving bijection from the set $\mathscr{R}$ of radical thick $\otimes$-ideals to the set $\mathscr{S}$ of open subsets of $\operatorname{Spc}^{cp}(\mathscr{K})^{\vee}$ induced by the assignment
$\mathscr{I}\mapsto\bigcup_{x\in\mathscr{I}}\operatorname{supp}^{cp}(x).$
In particular, every radical thick $\otimes$-ideal of $\mathscr{K}$ is an intersection of completely prime thick $\otimes$-ideals of $\mathscr{K}$.
@@
This essentially follows from [Kra24, Section 6]; we summarize here. For two $x,y\in\mathscr{K}$, write $x\sim y$ if and only if $\langle x\rangle=\langle y\rangle$, where $\langle x\rangle$ denotes the radical thick $\otimes$-ideal generated by $x$. Then $\sim$ is an equivalence relation, and $L(\mathscr{K},\otimes):=\operatorname{Ob}(\mathscr{K})/\sim$ is partially ordered via inclusion. [Kra24, Proposition 14] asserts that $L(\mathscr{K},\otimes)$ is a distributive lattice, and its ideal lattice $\operatorname{Id}(L(\mathscr{K},\otimes))$ identifies with the lattice of radical thick $\otimes$-ideals of $\mathscr{K}$. The bijection of the first statement now follows from [Kra24, Corollary 10] if the prime ideals of the lattice $L(\mathscr{K},\otimes)$ are precisely the complete primes of $\mathscr{K}$, i.e. we have an equality of sets $\operatorname{Spc}^{cp}(\mathscr{K})=\operatorname{Spc}(L(\mathscr{K},\otimes))$. Indeed, this follows from the identity $\langle x\rangle\wedge\langle y\rangle=\langle x\otimes y\rangle$ of [Kra24, Lemma 13], so $\langle x\rangle$ is a completely prime thick $\otimes$-ideal of $\mathscr{K}$ if and only if it is prime as an element of $L(\mathscr{K},\otimes)$.
The latter statement follows from the inverse of the map, which sends an open set $U$ to $\{x\in L(\mathscr{K},\otimes)\mid\operatorname{supp}^{cp}(x)\subseteq U\}$, which is nothing more than the intersection of all primes of $L(\mathscr{K},\otimes)$, i.e. complete primes, not contained in $U$ (c.f. [Bal05, Lemma 4.8]). ∎
@@
One must be careful with the usage of Thomason subsets here, since it is not clear if every open $U^{cp}(x)$ of $\operatorname{Spc}^{cp}(\mathscr{K})$ is quasi-compact - however, this indeed holds if every prime is complete. In fact, when $\operatorname{Spc}(\mathscr{K})=\operatorname{Spc}^{cp}(\mathscr{K})$, we informally note that nearly every topological property proven in [Bal05, Section 2] holds, essentially due to the crucial prime lifting lemma Proposition 2.4 holding for complete primes.